# PyAtoms: An interactive tool for rapidly simulating atomic scanning tunneling microscopy images of 2D materials, moiré systems and superlattices

### v2

@@ Abstract
[…]
1 Introduction
@@
Scanning tunneling microscopy (STM) [1] is a powerful technique for investigating the atomic-scale electronic and structural properties of materials [2]. In STM, an atomically sharp tip is rastered across a conductive sample surface while recording the current of electrons tunneling from the tip to the sample, or vice versa. The tunneling current is proportional to the energy-integrated local density of states (LDOS), and can thus provide valuable insights into the spatially-dependent electronic properties of novel quantum materials such as high-temperature superconductors [3, 4], charge- or bond-density wave systems [5, 6, 7], and two-dimensional (2D) van der Waals structures and moiré superlattices [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
@@
A drawback of STM measurements is that they are notoriously time-consuming: Depending on the particular imaging mode (topography or spectroscopy), the time to complete a measurement can take minutes to several days. In a STM topography measurement, the STM tip rapidly scans the surface (typical scan rate of $\sim 1$ Hz) at a single tunneling energy while recording the tunnel current line-by-line; in a STM spectroscopy measurement, the tunnel current and/or the differential conductance, $dI/dV_{b}\propto$ LDOS($eV_{b}$), (where $I$ is the tunnel current, $V_{b}$ is the voltage bias, and $e$ is the elementary charge), is slowly measured pixel-by-pixel at one or several tunneling energies. The choice of spatial resolution is especially important for spectroscopic measurements, where the Fourier-transformed STM (FT-STM) image can display intricate electronic scattering patterns – called quasi-particle interference (QPI) patterns – that are directly related to constant energy cuts of the momentum-resolved electronic band structure [4, 19]. The significant time commitment of STM imaging in real- or reciprocal-space thus makes it absolutely crucial for researchers to somehow determine the suitable imaging parameters prior to engaging a long-term measurement. This would require a tool that could easily simulate STM data in real- and Fourier-space for a variety of quantum material families while also providing real-time estimation of the measurement time.
@@
To address this need, we created PyAtoms: an open-source, Python-based graphical user interface (GUI) equipped with robust capabilities for quickly and easily simulating STM images of several types of 2D van der Waals quantum materials, including graphene, transition metal dichalcogenides, and systems with superlattices such as charge density waves and multi-layered moiré lattices. PyAtoms relies on a simplified model of an ideal atomic lattice, enabling the simulation of STM images for several families of quantum materials and allowing users to tune lattice strain and sublattice asymmetries. A notable strength of PyAtoms is its unique ability to rapidly simulate even complex STM and FT-STM moiré patterns in real-time. Users can use PyAtoms to quickly determine the local twist and/or lattice mismatch in their experimental data. Moreover, owing to its user-friendly GUI, we have found that PyAtoms is a valuable educational tool for teaching concepts related to periodic structures in real and reciprocal space. PyAtoms is thus a versatile and accessible software tool that can be utilized for novice and advanced 2D material researchers and educators.
@@
To our knowledge, no other software exists that combines PyAtoms’ level of 2D quantum material STM image simulation with a real-time GUI. Gwyddion [20], a versatile scanning probe microscopy (SPM) data analysis software, contains a module for simulating well-known surfaces (square, triangular, silicon $7\times 7$), but it lacks real-time simulation of multi-layer systems or for tuning strain or sublattice strength. Stephens and Hollen have shared a Matlab GUI for simulating moiré patterns using dot lattices that simultaneously display the Fourier transform [21]. de Jong has shared a very impressive lattice and moiré atomic simulator that allows for tuning the symmetry, strain, and sub-lattice, but it lacks a real-time GUI [22]. Finally, we note that, while PyAtoms is based on idealized atomic lattices, the National Institute of Standards and Technology (NIST) Joint Automated Repository for Various Integrated Simulations (JARVIS) database hosts a robust library of STM images calculated from density functional theory (https://jarvis.nist.gov/jarvisstm) [23].
@@
PyAtoms is built on Python 3 and utilizes widely used libraries such as NumPy [24], SciPy [25], Matplotlib [26] and PyQt5 (https://www.riverbankcomputing.com/software/pyqt). These libraries are readily available through common platforms such as Anaconda (https://www.anaconda.com). PyAtoms is hosted on Github (https://github.com/asariprado/PyAtoms) and is regularly updated. PyAtoms runs on both Windows and MacOS systems, however its compatibility with Linux systems has not been tested.
@@
The article is organized as follows: We first describe the theoretical underpinnings of PyAtoms. We describe how it generates single lattices of various geometries (square, triangular); how we approximate images of moiré and superlattice systems; and how strain and low-pass filtering affect the generated lattice images. We then describe the PyAtoms graphical interface before concluding with its applications in research and education.
2 Methods
2.1 The ideal atomic lattice
@@
The topographic intensity, $T$, of an ideal crystal at position $\mathbf{r}=(x,y)$ can be expressed as a Fourier sum [27, 28],
@@
where $\mathbf{g}_{k}$ corresponds to the $k^{th}$ reciprocal lattice vector corresponding to a Bragg reflection, and $A_{k}$ are the corresponding Fourier amplitudes for the $k^{th}$ mode which, in general, may be complex [27], and $\mathbf{r}_{0}$ is the location of the maxima of a reference atom in the image. For the sake of brevity in the following, we assume this reference atom is located at the origin $\mathbf{r}_{0}=(0,0)$. For a real image ($A_{k}=A_{-k}$), the image intensity is thus given by:
@@
Here, $N_{i}$ are constants ensuring the normalization of the images, such that $0\leq T(\mathbf{r})\leq 1$. Including higher harmonics of $\mathbf{g}_{k}$ (i.e. more Bragg peaks) in the sum in Eqs. 5, 6 will produce atoms that are more localized in space, approaching Dirac delta functions as the number of harmonics increases [21, 22].
@@
2.2 Honeycomb crystal structures
The honeycomb crystal, such as that in graphene, consists of a triangular Bravais lattice with a two-atom basis. It can be easily pictured as two interpenetrating triangular Bravais lattices, labeled $A$ and $B$ (Fig. 1(a)). Ignoring normalization for the moment, the $A$ sublattice centered at the origin, $\mathbf{r}_{0}=(0,0)$, and with amplitude $\alpha$ is given by
@@ 2.2 Honeycomb crystal structures
The $B$ sublattice centered at the nearest-neighbor distance $\mathbf{s}=(a/\sqrt{3})\hat{y}$ with respect to the $A$ sublattice, and with amplitude $\beta$, is given by
$T_{B}=\beta\sum_{k}\exp[i\mathbf{g}^{tri}_{k}\cdot(\mathbf{r}+\mathbf{s})]=\beta e^{2\pi i/3}\sum_{k}\exp(i\mathbf{g}^{tri}_{k}\cdot\mathbf{r}),$ (10)
@@
Figure 1: Generalized triangular lattices. (a) PyAtoms simulated image of graphene ($a=2.46$) created with equal amplitude sublattices ($\alpha=\beta=0.5$). (b) Vertical line scan of the image in (a) connecting consecutive hollow sites. (c) Series of simulated generalized honeycomb crystals with variable amplitudes on each sublattice. (d) Experimental STM image ($V_{b}=-400$ mV, $I_{t}=250$ pA, $T=$ 4.7 K) of graphite displaying sublattice asymmetry. (e) PyAtoms simulation of (d) ($\alpha=0.45,\beta=0.55$) that faithfully reproduces the sublattice asymmetry.
@@
Figure 1 displays several simulated STM images of a honeycomb crystal centered at a hollow (H) site. Ideal graphene (lattice constant $a=0.246$ nm) is simulated when the amplitudes of the $A$ and $B$ sublattices are equal (Fig. 1(a)). A vertical line cut of the simulated image (Fig. 1(b)) shows that the honeycomb crystal (solid black line) is composed of two triangular lattices, with atoms from each lattice centered at $A$ (solid red line) and $B$ (solid blue line). Modifying the relative amplitudes of the two sublattices $(\alpha,\beta)$ can create simulated STM images of more general honeycomb structures (Fig. 1(c)). This can be used to model STM images of transition metal dichalcogenides, graphene with broken sublattice symmetry, or experimental STM images recorded with an asymmetric tip. The relative sublattice strength $(\alpha,\beta)$ and the image origin ($A$, $B$, or hollow site) is easily implemented using PyAtoms’s GUI (Fig. 3(a)).
@@
As an example of PyAtoms’ robust capabilities, consider Fig. 1(d), which displays an experimental STM image of the surface of graphite. Typically, the surface of graphite appears in STM as a triangular lattice [1, 29]. The presence of the other sublattice is thus an artifact, most likely due to an asymmetric or doubled STM tip. This is easily simulated with PyAtoms in Fig. 1(e) by choosing sublattice parameters ($\alpha=0.45,\beta=0.55$).
2.3 Superlattices and multi-layered systems
@@
A superlattice is a larger periodic structure that is superimposed over an existing lattice. Superlattices can emerge from surface reconstructions [1], alloys and interfaces of materials [30], the presence of periodic lattice distortions and/or charge- or bond-density waves [5, 6, 31, 32], or, in moiré systems, from the combination of two or more crystal lattices [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. The moiré superlattice (wavevectors $\mathbf{g}_{M,i}$) can be understood as the real-space periodic analog of the beat-frequency in sound: Interference between two periodic lattices (labeled $A$ and $B$ with wavevectors $\mathbf{g}_{A,i}$ and $\mathbf{g}_{B,i}$) will create a moiré superlattice with a wavevector given by the difference, $\mathbf{g}_{M,i}=\mathbf{g}_{A,i}-\mathbf{g}_{B,i}$. The smaller the difference between the two lattices – either in magnitude or direction, i.e. twist – the smaller the moiré wavevector, and thus the longer the real-space wavelength, $\lambda_{M}\propto|\mathbf{g}_{M}|^{-1}$. The presence of this long-wavelength moiré periodic potential can have a dramatic effect on the electronic properties of 2D material stacks, from inducing superconductivity to correlated insulating behavior and magnetism [33, 34, 35, 36, 37].
@@
Figure 2: Moiré simulations of experimental data. (a) Experimental STM image of twisted graphene-graphite with a moiré twist of $\theta=1.9^{\circ}$. ($V_{b}=-400$ mV, $I_{t}=250$ pA, $T=$ 4.7 K). The height has been normalized to $[0,1]$ (b) Simulation using the “simple” moiré model with $\theta_{12}=1.9^{\circ},\eta=0.4$. (c) Same as (b) using the “log” moiré model with $\xi=0.85$. Note the compressed color scale. (d) Linecuts between the two AA moiré sites in (a)-(c). (e-f)Same as (a-c) after low-pass filtering ($\sigma=6$ pixels) which highlights the differences between the simulated moiré models. All images have a lateral size of 18.7 nm $\times$ 18.7 nm.
@@
Here, $T_{M}^{S}$ represents the moiré or other superlattice image, $T_{l}$ represents the individual constituent lattices, and $\eta$ is a parameter ($0\leq\eta\leq 1$) that weighs the relative strength of the sum of the lattices, $\sum_{l}T_{l}$, to their product, $\prod_{l}T_{l}$. Our toy model has notable strengths: (i) a single toy parameter, $\eta$, that is GUI-friendly; (ii) it creates real-space simulations with a wide tuneability in the image contrast to best compare with experimental data; and (iii) low-pass filtering of the resultant image agrees with experimental data (described below and Fig. 2f). We note that, with this model, the FT-STM image — by design — contains only first-order Bragg and superlattice peaks, which is typical of most experimental STM data [7, 9, 10, 38].
Tunneling or “Log” model: An alternative model proposed by Joucken et al [38] for simulating multi-layer moiré superlattices is rooted in the STM constant-current tunneling process and additionally takes into account the inter-atomic layer spacing, $d$, and the out-of-plane exponential decay length, $\lambda$. Here, the STM tunnel current is given by,
@@
where $z(x,y)$ is the STM tip height at lateral position $(x,y)$ and $T_{l}$ ($l=1,2,...)$ describes the LDOS in the $l$th atomic layer from the surface. In an STM constant-current measurement, the tip height, $z(x,y)$, changes dynamically to maintain a constant tunnel current, $I(x,y,z)$, through a feedback-loop. Setting the left-hand side of Eq. 14 constant and solving for the tip height, $z(x,y)$, yields the resultant superlattice STM topographic height image, $T^{L}_{M}=z(x,y)$, given by
@@
where, in the last line, we have combined the ratio of the inter-layer distance ($d$) to the decay length ($\lambda$) into a single, GUI-friendly parameter, $\xi=d/\lambda$: $\xi$ small corresponds to a short interlayer distance (or long atomic decay length); $\xi$ large corresponds to a large interlayer distance (or short atomic decay length). This model has its own notable strengths: (i) a single parameter, $\xi$, that is GUI-friendly; (ii) the relative intensity of constituent lattices can be tuned by adjusting $\xi$, and (iii) the Fourier-transformed image contains all orders of the superlattice peaks. However, the logarithm in Eq. 15, strongly affects the simulated image’s dynamic range which makes adjusting the contrast to match experimental data much more difficult. By allowing the user to toggle between the two models (“Simple” or “Log”), the user can determine the model and parameters to best match their experiment. Future versions of PyAtoms can easily allow even more models. However, we must stress that PyAtoms provides simulations of the ideal moiré lattices and cannot account for lattice relaxation [39].
Figure 2 contrasts the two models. Shown in Fig. 2(a) is an experimental STM image recorded on graphite at $T=$ 4.7 K. The presence of a moiré pattern indicates the delamination and twisting of the top-most graphene layer [40, 41]. The twist angle can be determined from the moiré wavelength using well-known geometric relations which yields $\theta_{12}\approx 1.9^{\circ}$. Alternatively, the parameter can be determined using PyAtoms, which can rapidly simulate the data as shown in Fig. 2(b,c) using either the “Simple” or “Log” model. The key differences between the two models can be seen by taking a linecut connecting the moiré AA sites (Fig. 2(d)). The “Simple” model (red line) has an amplitude similar to that in experiment (black line). The “Log” model (blue line), however, has a much narrower height range and, owing to Eq. 15, additionally features deep minima at the AA sites. Low-pass filtering the experimental and simulated images (Fig. 2(e)-(g)) shows that the “Simple” model more closely follows the experimental moiré image while the “Log” model yields an inverted filtered image owing to the deep minima at the AA sites.
2.4 Homogeneous strain

### v3

# PyAtoms: An interactive tool for simulating atomic scanning tunneling microscopy images of 2D materials, moiré systems and superlattices

@@ Abstract
We present PyAtoms, an interactive open-source software that rapidly simulates atomic-scale scanning tunneling microscopy (STM) and other scanning probe microscopy (SPM) images of two-dimensional (2D) layered materials, moiré systems, and superlattices. Rooted in a Fourier-space description of ideal atomic lattice images, PyAtoms is a Python-based graphical user interface (GUI) with robust capabilities for tuning lattice parameters (lattice constants, strain, number of layers, twist angles) and STM imaging parameters (pixels, scan size, scan angle) and provides time estimates for spectroscopic measurements. These capabilities allow users to efficiently plan time-consuming STM experiments. We provide an overview of PyAtoms’ current features, describe its underlying mathematical principles, and then demonstrate simulations of several 2D materials including graphene with variable sublattice asymmetry, twisted tri-layer graphene moiré systems, and several charge- and bond-density wave systems.
I Introduction
@@
Scanning tunneling microscopy (STM) binnig_scanning_1987 is a powerful technique for investigating the atomic-scale electronic and structural properties of materials moler_imaging_2017. In STM, an atomically sharp tip is rastered across a conductive sample surface while recording the current of electrons tunneling from the tip to the sample, or vice versa. The tunneling current is proportional to the energy-integrated local density of states (LDOS), and can thus provide valuable insights into the spatially-dependent electronic properties of novel quantum materials such as high-temperature superconductors Fischer_RevModPhys.79.353; hoffman_imaging_2002, charge- or bond-density wave systems Sacks_PhysRevB.57.13118; soumyanarayanan2013quantum; arguello_visualizing_2014; gutierrez_imaging_2016; gye2019topological; pasztor_holographic_2019, and two-dimensional (2D) van der Waals structures and moiré superlattices rutter2007scattering; xue2011scanning; decker_local_2011; zhang_interlayer_2017; kerelsky_maximized_2019; jiang_charge_2019; xie_spectroscopic_2019; li_imaging_2021; kim_imaging_2023; nuckolls_quantum_2023; slot2023quantum; ganguli2023visualization.
@@
A drawback of STM measurements is that they are notoriously time-consuming: Depending on the particular imaging mode (topography or spectroscopy), the time to complete a measurement can take minutes to several days. In a STM topography measurement, the STM tip rapidly scans the surface (typical scan rate of $\sim 1$ Hz) at a single tunneling energy while recording the tunnel current line-by-line; in a STM spectroscopy measurement, the tunnel current and/or the differential conductance, $dI_{t}/dV_{b}\propto$ LDOS($eV_{b}$), (where $I_{t}$ is the tunnel current, $V_{b}$ is the voltage bias, and $e$ is the elementary charge), is slowly measured pixel-by-pixel at one or several tunneling energies. The choice of spatial resolution is especially important for spectroscopic measurements, where the Fourier-transformed STM (FT-STM) image can display intricate electronic scattering patterns – called quasi-particle interference (QPI) patterns – that are directly related to constant energy cuts of the momentum-resolved electronic band structure hoffman_imaging_2002; avraham_quasiparticle_2018. The significant time commitment of STM imaging in real- or reciprocal-space thus makes it absolutely crucial for researchers to determine the suitable imaging parameters prior to engaging a long-term measurement. Recently, other scanning probe microscopy (SPM) techniques such as piezoresponse force microscopy (PFM) and torsional force microscopy (TFM) performed in ambient conditions have become widely used tools for rapid characterization and quality control of 2D moiré samplesmcgilly_visualization_2020; pendharkar_torsional_2024. These techniques share the need for careful selection of the real-space parameters, e.g. scan range, required to capture the essential features for analyzing the data in reciprocal-space. Additionally, multilayer moiré samples have the added difficulty of interpreting images which may have several or many intersecting length scales in the form of atomic moiré and supermoiré patterns. Efficient planning of STM/SPM measurements would require a tool that could easily simulate STM data in real- and Fourier-space for a variety of quantum material families while also providing real-time estimation of the measurement time.
@@
To address this need, we created PyAtoms: an open-source, Python-based graphical user interface (GUI) equipped with robust capabilities for quickly and easily simulating STM images of several types of 2D van der Waals quantum materials, including graphene, transition metal dichalcogenides, and systems with superlattices such as charge density waves and multi-layered moiré lattices. PyAtoms relies on a simplified model of an ideal atomic lattice, enabling the simulation of STM images for several quantum materials and allowing users to tune lattice strain and sublattice asymmetries. These capabilities allow PyAtoms to rapidly simulate even complex STM and FT-STM moiré patterns in real-time in order to, for instance, optimize scanning parameters before a time-consuming measurement or quickly determine local twist and/or lattice mismatch by side-by-side comparison with experimental data. PyAtoms is thus a versatile and accessible software tool for scanning probe and 2D material researchers.
@@
To our knowledge, no other software exists that combines PyAtoms’ level of 2D quantum material STM image simulation with a real-time GUI. Gwyddion Necas2012, a robust scanning probe microscopy (SPM) data analysis software, contains a module for simulating well-known surfaces (square, triangular, silicon $7\times 7$), but it lacks real-time simulation of multi-layer systems or for tuning strain or sublattice strength. Stephens and Hollen have shared a Matlab GUI for simulating moiré patterns using dot lattices that simultaneously display the Fourier transform stephens_moire_nodate. de Jong has shared a very impressive lattice and moiré atomic simulator that allows for tuning the symmetry, strain, and sublattice, but it lacks a real-time GUI de_jong_moire_2021. Finally, while PyAtoms is based on idealized atomic lattices, the National Institute of Standards and Technology (NIST) Joint Automated Repository for Various Integrated Simulations (JARVIS) database hosts a robust library of STM images calculated from density functional theory (https://jarvis.nist.gov/jarvisstm) choudhary_computational_2021.
@@
PyAtoms is built on Python 3 and utilizes widely used libraries such as NumPy harris_array_2020, SciPy virtanen_scipy_2020, Matplotlib hunter_matplotlib_2007 and PyQt5 (https://www.riverbankcomputing.com/software/pyqt). These libraries are readily available through common platforms such as Anaconda (https://www.anaconda.com). PyAtoms is hosted on Github (https://github.com/asariprado/PyAtoms) and is regularly updated. PyAtoms runs on Windows, MacOS, and Linux systems.
@@
The article is organized as follows: We first describe the theoretical underpinnings of PyAtoms. We describe how it generates single lattices of various geometries (square, triangular); how we approximate images of moiré and superlattice systems; and how strain and low-pass filtering affect the generated lattice images. We then describe the PyAtoms graphical interface before concluding with its applications in research.
II Methods
II.1 The ideal atomic lattice
@@
The topographic intensity, $T$, of an ideal crystal at position $\mathbf{r}=(x,y)$ can be expressed as a Fourier sum hytch_quantitative_1998; lawler_intra-unit-cell_2010; hamidian_detection_2016,
@@
where $\mathbf{g}_{k}$ corresponds to the $k^{th}$ reciprocal lattice vector corresponding to a Bragg reflection, and $A_{k}$ are the corresponding Fourier amplitudes for the $k^{th}$ mode which, in general, may be complex hytch_quantitative_1998, and $\mathbf{r}_{0}$ is the location of the maxima of a reference atom in the image. For the sake of brevity in the following, we assume this reference atom is located at the origin $\mathbf{r}_{0}=(0,0)$. For a real image ($A_{k}=A_{-k}$), the image intensity is thus given by:
@@
Here, $N_{i}$ are constants ensuring the normalization of the images, such that $0\leq T(\mathbf{r})\leq 1$. Including higher harmonics of $\mathbf{g}_{k}$ (i.e. more Bragg peaks) in the sum in Eqs. 5, 6 will produce atoms that are more localized in space, approaching Dirac delta functions as the number of harmonics increases stephens_moire_nodate; latychevskaia_moire_2019; de_jong_moire_2021.
@@
Figure 1: Generalized triangular lattices. (a) PyAtoms simulated image of graphene ($a=2.46$ Å) created with equal amplitude sublattices ($\alpha=\beta=0.5$). (b) Vertical line scan of the image in (a) connecting consecutive hollow sites. (c) Series of simulated generalized honeycomb crystals with variable amplitudes on each sublattice. (d) Experimental STM image ($V_{b}=-400$ mV, $I_{t}=250$ pA, $T=$ 4.7 K) of graphite displaying sublattice asymmetry. (e) PyAtoms simulation of (d) ($\alpha=0.45,\beta=0.55$) that faithfully reproduces the sublattice asymmetry.
II.2 Honeycomb crystal structures
The honeycomb crystal, such as graphene, consists of a triangular Bravais lattice with a two-atom basis. It can be easily pictured as two interpenetrating triangular Bravais lattices, labeled $A$ and $B$ (Fig. 1(a)). Ignoring normalization for the moment, the AA sublattice centered at the origin, $\mathbf{r}_{0}=(0,0)$, and with amplitude $\alpha$ is given by
@@ II.2 Honeycomb crystal structures
The $B$ sublattice centered at the nearest-neighbor distance $\mathbf{s}=(a/\sqrt{3})\hat{y}$ with respect to the AA sublattice, and with amplitude $\beta$, is given by
$T_{B}=\beta\sum_{k}\exp[i\mathbf{g}^{tri}_{k}\cdot(\mathbf{r}+\mathbf{s})]=\beta e^{2\pi i/3}\sum_{k}\exp(i\mathbf{g}^{tri}_{k}\cdot\mathbf{r}).$ (10)
@@
Figure 1 displays several simulated STM images of a honeycomb crystal centered at a hollow (H) site. Ideal graphene (lattice constant $a=0.246$ nm) is simulated when the amplitudes of the AA and $B$ sublattices are equal (Fig. 1(a)). A vertical line cut of the simulated image (Fig. 1(b)) shows that the honeycomb crystal (solid black line) is composed of two triangular lattices, with atoms from each lattice centered at $A$ (solid red line) and $B$ (solid blue line). Modifying the relative amplitudes of the two sublattices $(\alpha,\beta)$ can create simulated STM images of more general honeycomb structures (Fig. 1(c)). This can be used to model STM images of transition metal dichalcogenides, graphene with broken sublattice symmetry, or experimental STM images recorded with an asymmetric tip. The relative sublattice strength $(\alpha,\beta)$ and the image origin ($A$, $B$, or hollow site) is easily implemented using PyAtoms’s GUI (Fig. 3(a)).
@@
As an example of PyAtoms’ robust capabilities, consider Fig. 1(d), which displays an experimental STM image of the surface of graphite. Typically, the surface of graphite appears in STM as a triangular lattice binnig_scanning_1987; hembacher2003revealing. The presence of the other sublattice is thus an artifact, most likely due to an asymmetric or doubled STM tip. This is easily simulated with PyAtoms in Fig. 1(e) by choosing sublattice parameters ($\alpha=0.45,\beta=0.55$).
II.3 Superlattices and multi-layered systems
@@
A superlattice is a larger periodic structure that is superimposed over an existing lattice. Superlattices can emerge from surface reconstructions binnig_scanning_1987, alloys and interfaces of materials carpinelli_direct_1996, the presence of periodic lattice distortions and/or charge- or bond-density waves Sacks_PhysRevB.57.13118; arguello_visualizing_2014; pasztor_holographic_2019; Shunsuke_PhysRevLett.132.056401, or, in moiré systems, from the combination of two or more crystal lattices rutter2007scattering; xue2011scanning; decker_local_2011; zhang_interlayer_2017; kerelsky_maximized_2019; jiang_charge_2019; xie_spectroscopic_2019; li_imaging_2021; kim_imaging_2023; slot2023quantum. The moiré superlattice (wavevectors $\mathbf{g}_{A,i,B,j}$) can be understood as the real-space periodic analog of the beat-frequency in sound: Interference between two periodic lattices (labeled $A$ and $B$ with wavevectors $\mathbf{g}_{A,i}$ and $\mathbf{g}_{B,j}$, where $i,j$ index the Bragg wavevectors) will create a moiré superlattice with wavevectors given by the difference, $\mathbf{g}_{A,i,B,j}=\mathbf{g}_{A,i}-\mathbf{g}_{B,j}$. The smaller the difference between the two lattices – either in magnitude or direction, i.e. twist – the smaller the moiré wavevector, and thus the longer the real-space wavelength, $\lambda_{M}\propto|\mathbf{g}_{A,i,B,j}|^{-1}$. The presence of this long-wavelength moiré periodic potential can have a dramatic effect on the electronic properties of 2D material stacks, from inducing superconductivity to correlated insulating behavior and magnetism cao_correlated_2018; cao_unconventional_2018; sharpe_emergent_2019; balents2020superconductivity; mak2022semiconductor.
@@
Figure 2: Moiré simulations of experimental data. (a) Experimental STM image of twisted graphene-graphite with a moiré twist of $\theta=1.9^{\circ}$. ($V_{b}=-400$ mV, $I_{t}=250$ pA, $T=$ 4.7 K). The height has been normalized to $[0,1]$ . (b) Simulation using the “simple” moiré model with $\theta_{12}=1.9^{\circ},\eta=0.4$. (c) Same as (b) using the “log” moiré model with $\xi=0.85$. Note the compressed color scale. (d) Linecuts between the two $A$ moiré sites in (a)-(c). (e-f)Same as (a-c) after low-pass filtering ($\sigma=6$ pixels) which highlights the differences between the simulated moiré models. All images have a lateral size of 18.7 nm $\times$ 18.7 nm.
@@
Here, $T_{M}^{S}$ represents the moiré or other superlattice image, $T_{l}$ represents the individual constituent lattices, and $\eta$ is a parameter ($0\leq\eta\leq 1$) that weighs the relative strength of the sum of the lattices, $\sum_{l}T_{l}$, to their product, $\prod_{l}T_{l}$. Our toy model has notable strengths: (i) a single toy parameter, $\eta$, that is GUI-friendly; (ii) it creates real-space simulations with a wide tuneability in the image contrast to best compare with experimental data; and (iii) low-pass filtering of the resultant image agrees with experimental data (described below and Fig. 2f). Using this model, the FT-STM image — by design — contains only first-order Bragg and superlattice peaks, which is typical of most experimental STM data gutierrez_imaging_2016; xue2011scanning; decker_local_2011; joucken_fourier_2015.
Tunneling or “Log” model: An alternative model proposed by Joucken et al joucken_fourier_2015 for simulating multi-layer moiré superlattices is rooted in the STM constant-current tunneling process and additionally takes into account the inter-atomic layer spacing, $d$, and the out-of-plane exponential decay length, $\lambda$. Here, the STM tunnel current is given by,
@@
where $z(x,y)$ is the STM tip height at lateral position $(x,y)$ and $T_{l}$ (l=1,2,)l=1,2,...) describes the LDOS in the $l$th atomic layer from the surface. In an STM constant-current measurement, the tip height, $z(x,y)$, changes dynamically to maintain a constant tunnel current, $I(x,y,z)$, through a feedback-loop. Setting the left-hand side of Eq. 14 constant and solving for the tip height, $z(x,y)$, yields the resultant superlattice STM topographic height image, $T^{L}_{M}=z(x,y)$, given by
@@
where, in the last step, we have combined the ratio of the inter-layer distance ($d$) to the decay length ($\lambda$) into a single parameter, $\xi=d/\lambda$: $\xi$ small corresponds to a short interlayer distance (or long atomic decay length); $\xi$ large corresponds to a large interlayer distance (or short atomic decay length). This model has its own notable strengths: (i) a single parameter, $\xi$, that is GUI-friendly; (ii) the relative intensity of constituent lattices can be tuned by adjusting $\xi$, and (iii) the Fourier-transformed image contains all orders of the superlattice peaks. However, the logarithm in Eq. 15, strongly affects the simulated image’s dynamic range which makes adjusting the contrast to match experimental data much more difficult. By allowing the user to toggle between the two models (“Simple” or “Log”), the user can determine the model and parameters to best match their experiment.
Figure 2 contrasts the two models. Shown in Fig. 2(a) is an experimental STM image recorded on graphite at $T=$ 4.7 K. The presence of a moiré pattern indicates the delamination and twisting of the top-most graphene layer li_scanning_2009; li_observation_2010. The twist angle can be determined from the moiré wavelength using well-known geometric relations which yields $\theta_{12}\approx 1.9^{\circ}$. Alternatively, the parameter can be determined using PyAtoms, which can rapidly simulate the data as shown in Fig. 2(b,c) using either the “Simple” or “Log” model. The key differences between the two models can be seen by taking a linecut connecting the moiré AA sites (Fig. 2(d)). The “Simple” model (red line) has an amplitude similar to that in experiment (black line). The “Log” model (blue line), however, has a much narrower height range and, owing to Eq. 15, additionally features deep minima at the AA sites. Low-pass filtering the experimental and simulated images (Fig. 2(e)-(g)) shows that the “Simple” model more closely follows the experimental moiré image while the “Log” model yields an inverted filtered image owing to the deep minima at the AA sites.